# Characterizing the Complexity of Social Robot Navigation Scenarios

### v1

@@ IV-D Results
H1. We find a statistically significant correlation between Density and Success rate and Minimum Distance to agent ($\rho=-0.867,-0.735$, $p<0.001$ using Spearman’s r test), and a moderate correlation between Density and average Time and Path Irregularity ($\rho=0.499,0.463$, $p<0.001$). We also find a correlation between increasing width and method Success ($\rho=0.593$, $p<0.001$), and a correlation between width and Minimum Distance ($\rho=0.285$, $p<0.05$). We do not find any statistically significant correlation with average Time or Path Irregularity. Regarding Policy Mixture and Directionality, we find strong and moderate correlations respectively between higher intensity and Success ($\rho=-0.715,-0.462$, $p<0.002$), and a moderate correlation between Directionality and average Time ($\rho=-0.314$, $p<0.05$). We see no statistically significant correlation for Minimum Distance, and Path Irregularity in either case. Thus, we find that while each factor affects at least one metric, Density and Environment Geometry appear to have the strongest correlation, giving partial support for H1.
@@ IV-E Analysis
RGL achieves higher Success than the MPC and MPPI methods in nearly all experiments, particularly in high density scenarios. Its Success rate performance nearly mirrors ORCA (its training Policy Mixture), however with consistently higher Minimum Distances to humans. This indicates it learned similar collision avoidance behavior to ORCA, while also learning to better respect agents’ personal space.
MPC-CV, MPC-SGAN, and MPPI-SGAN performance scales poorly with moderate to high crowd density. We do observe that MPPI-SGAN consistently has higher Success rates than MPC-SGAN, showing that a stronger controller implementation does indeed lead to better performance (albeit with slightly lower Minimum Distances). The steep decrease of all three methods, however, indicates that having an inaccurate prediction model becomes increasingly problematic as the number of agent trajectories predicted increases, even with a more robust controller.
@@
We do however see that the MPC methods maintain the highest Minimum Distance, while ORCA, RP, and CV have the lowest Minimum Distances. Additionally, the MPC methods, with the exception of the Environment Width experiments, maintain comparable Path Irregularity to ORCA, SFM, and RGL, showing they can maintain distance with comparably smooth paths. Thus we see that while the reactive and CV methods (with the exception of SFM) achieve efficient, collision-free navigation through other agents’ cooperation, they are worse at respecting the personal space of others compared to predictive methods.
We obtained some evidence that increased heterogeneity of scenarios is not by itself an indicator of increased Complexity. Qualitatively, we observed that passing scenarios are generally easier to navigate than crossing, and SFM agents are much more subservient than ORCA. An implicit hypothesis in our experimental ordering was that a mixture of two Directionalities or Policy Mixtures would be more Complex to navigate than either individually. We instead see improvement in the ORCA/SFM and Passing/Crossing scenarios compared to ORCA and Crossing only, which suggests the combinations are actually less Complex.
@@ V Discussion
Conducting high-Complexity evaluations. Evaluations in the literature are typically conducted under low-Density settings (Table I). Our experiments demonstrate that most algorithms handle those well but experience significant performance reductions as Density increases. Similarly, while many works contain passing-only experiments, in our study, Passing has the highest Success rates and Minimum Distances to agents, indicating it is the easiest to navigate. Additionally, while many works consider Medium-to-Large rooms, these appear to be the simplest to navigate (Fig. 3). Finally, nearly all evaluations in recent literature focus on scenarios with exclusively cooperative agents, which our experiments showed are easiest. These practices hinder the extraction of helpful insights; for example, had we not manipulated the Complexity factors towards the upper extremes, the severe performance drops at high density experienced by MPC-CV, MPC-SGAN, and MPPI-SGAN would not have been identified. These observations suggest that the social navigation community should shift to studying scenarios of higher Complexity to move beyond the frontier. Specifically, more evaluations in more geometrically-constrained, highly-mixed and high-traffic environments, similar to many context-rich public domains should be conducted.
Handling test-time distribution. The high Success rates of CV and RP suggest that in lower-Complexity regimes, it might be sufficient to use non-reactive classical planning techniques. As the Complexity increases, all algorithms experience a steep decline; the magnitude of the decline is related to how far out of distribution an algorithm operates. Our evaluation (Fig. 3) captures the sensitivity of data-driven approaches to their training distribution. For instance, we see that RGL, and MPC/MPPI-SGAN experience a substantial performance decline as the test-time Density moves away from the training/tuning Density ($0.05,0.10$ $agents/m^{2}$). In practice, data-driven approaches will often face distribution shift at deployment, which motivates the development of algorithms for overcoming distribution shifts to be an important direction for future research [30, 73]. Alternative approaches include techniques for proactive, expressive motion generation [34, 3, 66] or techniques that embrace graceful touch [43], similar to how human crowds resolve close interactions in train stations or airports [19].
@@
Simulating complex settings. While realistically simulating pedestrian-robot interactions is challenging [14, 36], we view it as an essential tool for testing high-Complexity settings that are difficult to safely replicate in the real world. This requires revisiting conventional assumptions, such as that humans are non-reactive to the robot [38, 29, 68, 32, 9, 8], which is unrealistic since the robots used (see Table I) have a large enough footprint to be observed. Thus, simulation should focus on visible robot settings, leveraging metrics and considerations of users’ perceptions [34, 53, 66] to compare algorithms’ social performance.
Limitations. Our analysis is a step towards understanding the Complexity of social navigation scenarios; an important next step would be to validate our observations with real-world experiments. While our evaluation involves standard metrics [36, 15], it lacks a user-centered perspective. Future work will incorporate metrics mapping robot behavior to users’ impressions [11, 34, 66]. We tested algorithms for which we found open-source implementations but additional algorithms could be included. We strived for a fair comparison by tuning all algorithms wrt the same criteria. While tuning can always be improved, the key takeaway remains: there is a steep drop in performance as Complexity increases. Finally, our analysis only explored the impact of contextual factors on robot performance; future work will look into the robot-related factors of kinematics, sensors, and viewpoints.

### v2

@@ IV-D Results
H1. We find Density has a strong correlation with Success Rate and Minimum Distance to agent ($\rho=-0.878,-0.760$, $p<0.001$ using Spearman’s r test), and a moderate correlation with Average Time ($\rho=0.488$, $p<0.001$). We also find a strong correlation between increasing Width and Success Rate ($\rho=0.641$, $p<0.001$), and a moderate correlation between Width and Minimum Distance ($\rho=0.420$, $p<0.01$). We do not find any statistically significant correlation with average Time or Path Irregularity. Regarding Policy Mixture and Directionality, we find strong and moderate correlations respectively between intensity and Success Rate ($\rho=-0.778,-0.524$, $p<0.002$). We see no statistically significant correlation for Minimum Distance and Time in either case, and no statistically significant correlations for Path Irregularity. Thus, we find that while each factor affects at least one metric, Density and Environment Geometry appear to have the strongest correlation, giving partial support for H1.
@@ IV-E Analysis
Despite being trained on only 5 ORCA agent circle crossings, RGL achieves higher Success than the MPC and MPPI methods in nearly all experiments, particularly in high Density scenarios. Its Success Rate nearly mirrors ORCA (its training Policy Mixture), however with consistently higher Minimum Distances to humans. This indicates it learned similar collision avoidance behavior to ORCA, while also learning to better respect agents’ personal space.
MPC-CV, MPC-SGAN, and MPPI-SGAN performance scales poorly with moderate to high Density, although we do see MPPI-SGAN scales better than MPC-SGAN, showing that a stronger controller implementation does indeed lead to better performance (albeit with slightly lower Minimum Distances). The steep decrease for all three methods, however, indicates that having an inaccurate prediction model becomes increasingly problematic as the number of agent trajectories predicted increases, even with a more robust controller.
@@
We do however see that the MPC methods maintain the highest Minimum Distance, while ORCA, RP, and CV have the lowest Minimum Distances. Additionally, the MPC methods, with the exception of the Environment Width experiments, maintain comparable Path Irregularity to ORCA, SFM, and RGL, showing they can maintain distance with comparably smooth paths. Thus we see that while the reactive and CV methods (with the exception of SFM) achieve efficient, collision-free navigation among cooperative agents, they are worse at respecting the personal space of others compared to predictive methods.
We obtained evidence that increased heterogeneity of scenarios is not by itself an indicator of increased Complexity. Qualitatively, we observed that passing scenarios are generally easier to navigate than crossing, and SFM agents are much more subservient than ORCA. An implicit hypothesis in our experimental ordering was that a mixture of two Directionalities or Policy Mixtures would be more Complex to navigate than either individually. We instead see improvement in the ORCA/SFM and Passing/Crossing scenarios compared to ORCA and Crossing only, which suggests the combinations are actually less Complex.
@@ V Discussion
Conducting high-Complexity evaluations. Evaluations in the literature are often conducted under low-Density settings (Table I). Our experiments demonstrate that most algorithms handle those well but experience significant performance reductions as Density increases. Similarly, while many works contain passing-only experiments, in our study, Passing has the highest Success rates and Minimum Distances to agents, indicating it is the easiest to navigate. Additionally, while many works consider Medium-to-Large rooms, these appear to be the simplest to navigate (Fig. 3). Finally, nearly all evaluations in recent literature focus on scenarios with exclusively cooperative agents, which our experiments show are easiest. These practices hinder the extraction of helpful insights; for example, had we not manipulated the Complexity factors towards the upper extremes, the severe performance drops at high Density experienced by MPC-CV, MPC-SGAN, and MPPI-SGAN would not have been identified. These observations suggest that the social navigation community should shift towards studying scenarios of higher Complexity to move beyond the frontier. Specifically, more evaluations in more geometrically-constrained, highly-mixed and high-traffic environments, similar to many context-rich public domains should be conducted.
Handling test-time distribution. The high Success rates of CV and RP suggest that in lower-Complexity regimes, it might be sufficient to use non-reactive classical planning techniques. As the Complexity increases, all algorithms experience a steep performance decline, the magnitude of which serves to indicate how far out of distribution an algorithm operates. Our evaluation (Fig. 3) captures the sensitivity of data-driven approaches to their training distribution. For instance, we see that RGL, and MPC/MPPI-SGAN experience a substantial performance decline as the test-time Density moves away from the training/tuning Density ($0.05,0.10$ $agents/m^{2}$). In practice, data-driven approaches will often face distribution shift at deployment, leading to unexpected or unsafe behavior, as demonstrated by the performance drop experienced by RGL and the MPC approaches in scenarios of different Complexity than their training data. Thus, to safely and effectively deploy social navigation algorithms, it will be important to rigorously characterize the scenarios in which they can be expected to perform effectively, and which scenarios are out of distribution and may cause performance degradation.
@@
Simulating complex settings. While realistically simulating pedestrian-robot interactions is challenging [29, 1], we view it as an essential tool for testing high-Complexity settings that are difficult to safely replicate in the real world. This requires revisiting conventional assumptions, such as that humans are non-reactive to the robot [43, 13, 58, 42, 36, 64], which is unrealistic since the robots used (see Table I) generally have a large enough footprint to be observed. Thus, simulation should focus on visible robot settings, leveraging metrics and considerations of users’ perceptions [27, 28, 26] to compare algorithms’ social performance.
Limitations. Our analysis is a step towards understanding the Complexity of social navigation scenarios; an important next step would be to validate our observations with real-world experiments. While our evaluation involves standard metrics [1, 2], it lacks a user-centered perspective. Future work will incorporate metrics mapping robot behavior to users’ impressions [69, 27, 26]. We tested algorithms for which we found open-source implementations but additional algorithms could be included. We strived for a fair comparison by tuning all algorithms w.r.t. the same criteria. While tuning can always be improved, the key takeaway remains: there is a steep drop in performance as Complexity increases. Additionally, we acknowledge that our simulator makes the unrealistic assumptions of perfect sensing and constant-sized circular agents, and these could be improved in future implementations. Furthermore, our results are affected by our choice of ORCA and SFM to simulate cooperative agents, where other algorithms may have led to different results. While these assumptions reduce the realism of our experiments, they allow us to minimize the effect of extraneous factors on the relationship between the Complexity factors and performance. Future work will investigate the robot-related factors of kinematics, sensors, and viewpoints, and can explore the effects of additional real-world considerations.